A sample of chlorine gas occupies a volume of 946 mL at a pressure of 726 mmHg. What is the pressure of the gas (in mmHg) if the volume is reduced at constant temperature to 154 mL?
1
Expert's answer
2016-03-24T11:32:03-0400
P*V = P1*V1 (T=const)
P1 = P*V/V1
P1 = 726*0.946 / 0.154 = 594578028 Pa = 4459.714 mmHg
Answer
P1 = 4459.714 mmHg